Coeffecient of Variation

Monarch King Size Beds Ltd. is evaluating a new promotional campaign that could increase sales. Possible outcomes and probabilities of the outcomes are shown below.  

 

  Possible OutcomesAdditional
Sales in Units
Probabilities
  Ineffective campaign
30

0.30
  Normal response
110

0.50
  Extremely effective
130

0.20

 

Compute the coefficient of variation. (Round the final answer to 2 decimal places.)

  

Coefficient of variation           
Answer #1

Expected value=Respective sales*Respective probability

=(0.3*30)+(0.5*110)+(0.2*130)=90

probabilitySalesprobability*(Sales-Expected Sales)^2
0.3300.3*(30-90)^2=1080
0.51100.5*(110-90)^2=200
0.21300.2*(130-90)^2=320


Total=1600

Standard deviation=[Total probability*(Sales-Expected Sales)^2/Total probability]^(1/2)

=40

Coefficient of variation=Standard deviation/Expected value

=40/90

=0.444(Approx).


answered by: biz.fan